Could I get some help with graphing this equation: 2x + y = -8? Thank You
Answer by AnlytcPhil(1810)

Could I get some help with graphing this equation: 2x + y = -8? 
Thank You

1.  Find the x-coordinate of the x-intercept by substituting 0 for y and
    solving for x.  The y-coordinate of the x-intercept is 0.

    2x + y = -8
    2x + 0 = -8
        2x = -8
         x = -8/2
         x = -4

    So the x-intercept is (-4,0)

2.  Find the y-coordinate of the y-intercept by substituting 0 for x and
    solving for y.  The x-coordinate of the y-intercept is 0.

    2x + y = -8
    2(0) + y = -8
       0 + y = -8
           y = -8
 
    So the x-intercept is (0,-8)
   
3. Plot these two points (-4,0) is on the x-axis 4 units left of the
   origin and (0,-8) is on the y-axis 8 units below the origin. 

           
    
Then draw a line through those two points:

          

Edwin